A refrigerator removes heat from the freezing compartment at the rate of 20kj and ejects 24 km into a rookper cycle. How much work is required in each cycle?
Answer #1

Solution)

We know, Work done,

W = Q2 - Q1 / n

Where, n= 1 - Q1 / Q2

Efficiency, n= 1 - 20 / 24

So, n = 0.16

Now,

W = Q2 - Q1 / n

Hence,

W = 24 - 20 / 0.16

W = 24 k J

===========